OpenStudy (shamil98):

What are the equations of the asymptotes of this parabola? x^2/25 - y^2/4 =1

OpenStudy (alekos):

This is not a parabola, its an hyperbola. For the general eqtn x^2/a^2 - y^2/b^2 =1 the asymptotes are y = (-b/a)x & (b/a)x

OpenStudy (anonymous):

h=0, k=0, a=5, b=2 slopes of asymptotes: m₁= b/a=2/5 m₂= -b/a= -2/5 The asymptotes pass through the center of the hyperbola. Use the center point to determine the y-intercepts of the asymptotes. center (h,k) = (0,0) y-intercept₁ = k - m₁h = 0 - (2/5)(0) = 0 y-intercept₂ = 0 - (-2/5)(0) = 0 asymptote₁: y = (2/5)x asymptote₂: y = (-2/5)x